# HyPINO: Multi-Physics Neural Operators via HyperPINNs and the Method of Manufactured Solutions

### v1

@@ 4.3 Evaluation
We evaluate HyPINO and baseline models on seven standard PDE benchmarks from the PINN literature: (i) HT - 1D heat equation Lu et al. (2021b), (ii) HZ - 2D Helmholtz equation Bischof and Kraus (2025), (iii) HZ-G - Helmholtz on an irregular geometry Hao et al. (2023b), (iv) PS-C - Poisson with four circular interior boundaries Hao et al. (2023b), (v) PS-L - Poisson on an L-shaped domain Lu et al. (2021b), (vi) PS-G - Poisson with a Gaussian vorticity field Herde et al. (2024), and (vii) WV - 1D wave equation Hao et al. (2023b). The exact problem statements and visualizations of the corresponding parameterizations are provided in Appendix B.
Table 1: Model performance across seven PDE benchmarks. Each cell shows mean squared error (MSE) / symmetric mean absolute percentage error (SMAPE) Makridakis (1993). Lower is better.
@@
Table 1 further highlights the advantages of our proposed iterative refinement approach. After three refinement iterations (HyPINO3), we observe substantial reductions in prediction error across all but one benchmark. Notably, the MSE for PS-C and PS-G decreases by more than one order of magnitude, and an even larger improvement is observed for HT (almost two orders of magnitude). With ten refinement iterations (HyPINO10), our model achieves state-of-the-art performance on all but two evaluated benchmarks, outperforming the best baseline models by factors ranging from 2.1 (on HZ against Poseidon) to 173 (HT against PINO). Table 2 shows that iterative refinement leads to a continuously more accurate prediction on the challenging WV benchmark, with the model being able to extend the undulating shape continuously further from the initial condition across the time dimension.
@@ 4.4 Finetuning
The parameters $\theta^{\star}$ produced by HyPINO can be used to initialize PINNs for subsequent fine-tuning on specific PDE instances. We compare the convergence behavior of HyPINO-initialized PINNs with those initialized randomly and with Reptile meta-learning Nichol et al. (2018), where Reptile was trained on our synthetic dataset using 10,000 outer- and 1,000 inner-loop cycles. We also evaluate ensembles generated with HyPINO3 and HyPINO10 against ensembles of equal size and architecture initialized with random weights or Reptile.
@@ A.1 Neural Operator Architecture
We build on the HyperPINN framework de Avila Belbute-Peres et al. [2021] and design a neural operator based on a hypernetwork that generates the weights $\theta^{\star}$ of a Physics-Informed Neural Network (PINN) $u_{\theta}$, conditioned on a given PDE instance. Specifically, the hypernetwork learns a mapping
@@
Each grid-valued input is first passed through a Fourier feature mapping Tancik et al. [2020], which augments the input with sinusoidal encodings using five exponentially increasing bands $\text{frequencies}=0.1\cdot 2^{i},\quad i\in\{0,1,2,3,4\}$. This enhances the network’s ability to represent high-frequency content and reduces spectral bias. The Fourier mapping layer is followed by two convolutional layers with kernel size three and strides of two. For the boundary location grids $M$ (c.f. Section 3.1, we compute four embeddings: $z_{D}^{1},\;z_{D}^{2},\;z_{N}^{1},\;z_{N}^{2}$. For the boundary value grids $V$, we compute $z_{g}$ (Dirichlet values) and $z_{h}$ (Neumann values). The source term yields the embedding $z_{f}$.
@@
The grid embedding $z_{G}$ is processed by a sequence of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$. Denoting by $z^{(i)}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ the output of block $\mathcal{SW}_{i}$ and $z^{(0)}=z_{G}$, we interleave each block with a FiLM modulation Perez et al. [2017] conditioned on the coefficient embeddings $z_{C}$. Concretely, we define
@@
Inspired by Swin Transformer UNet architectures Cao et al. [2021], Fan et al. [2022], we retain all intermediate latent representations from the Swin blocks $\{z^{(i)}\}_{i=1}^{K}$ to keep information at various semantic levels.
@@
To aggregate spatial information into a compact latent representation suitable for parameterizing the target PINN, we perform Multi-Head Attention Pooling Lee et al. [2019], Zhai et al. [2022] across the flattened outputs of each Swin Transformer block. Specifically, let $z_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ denote the output of the $i$-th FiLM-modulated Swin block. We reshape it into a sequence of tokens $kv_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$, which serve as both keys and values in the attention mechanism.
@@
We define the architecture of the target PINN as MLP with Fourier feature mapping Tancik et al. [2020] and multiplicative skip connections Wang et al. [2021a]. Fourier encodings provide spectral expressivity for modeling high-frequency components Wang et al. [2021b], while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by masking some layers.
@@
Following Wang et al. [2021a], the encoded input is projected through three parallel transformations:
@@ A.2.5 Sampling Physical Domains
We employ a randomized sampling procedure based on Constructive Solid Geometry (CSG) Lu et al. [2021c] to generate complex and diverse domains.
@@ B.1 Heat 1D (HT)
This benchmark problem is adapted from DeepXDE Lu et al. [2021b]. Figure 5 shows the parametrization of the different PDE components.
@@ B.2 Helmholtz 2D (HZ)
This benchmark problem is adapted from DeepXDE Lu et al. [2021b]. Figure 6 shows the parametrization of the different PDE components.
@@ B.3 Helmholtz 2D - Complex Geometry (HZ-G)
This benchmark problem is adapted from PINNacle Hao et al. [2023b]. Figure 7 shows the parametrization of the different PDE components.
@@ B.4 Poisson 2D - Circles (PS-C)
This benchmark problem is adapted from PINNacle Hao et al. [2023b]. Figure 8 shows the parametrization of the different PDE components.
@@ B.5 Poisson 2D - L-Domain (PS-L)
This benchmark problem is adapted from DeepXDE Lu et al. [2021b]. Figure 9 shows the parametrization of the different PDE components.
@@ B.6 Poisson 2D - Gauss (PS-G)
We select a sample from the dataset introduced in Herde et al. [2024]. Figure 10 shows the parametrization of the different PDE components.
@@ B.7 Wave 1D (WV)
This benchmark problem is adapted from PINNacle Hao et al. [2023b]. Figure 11 shows the parametrization of the different PDE components.
@@ B.8 Baseline Models
UNet Ronneberger et al. [2015]. A convolutional encoder–decoder network that shares the same input encoding architecture as HyPINO, but replaces the transformer-based decoder with a purely convolutional upsampling stack that directly outputs a solution grid of shape $(224\times 224)$, matching the resolution of the input tensors. It is trained exclusively on supervised PDEs with analytical solutions, using a batch size of 128, an initial learning rate of $10^{-4}$, and for 30,000 training batches. The UNet has a total parameter count of 62M.
Poseidon Herde et al. [2024]. A large pretrained operator network with approximately 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the input dimensionality to 5 to accept all grid-based inputs. Additionally, the lead-time-conditioned layer normalization layers—originally designed to condition on a 1D time input—are modified to condition on the 5D vector of differential operator coefficients. Poseidon is fine-tuned exclusively on supervised samples, using the same training setup as the UNet (30,000 batches, batch size 128, initial learning rate $10^{-4}$).
PINOLi et al. [2024]. A Fourier neural operatorLi et al. [2020] architecture with 33M parameters, trained with joint physics-informed and supervised losses computed in Fourier space. We adapt the model to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It follows the same hybrid supervision and training curriculum as HyPINO, including physics-informed losses, and is trained for 30,000 batches with a batch size of 128 and an initial learning rate of $10^{-4}$.
@@ B.10 Finetuning
We investigate the utility of HyPINO-generated PINN parameters $\theta^{\star}$ as a prior for rapid adaptation to specific PDE instances. We compare three initialization strategies: (i) HyPINO-initialized PINNs, (ii) randomly initialized PINNs, and (iii) PINNs initialized via Reptile meta-learning Nichol et al. [2018]. For Reptile, we use 10,000 outer-loop steps on our synthetic dataset and 1,000 inner-loop updates per sample with an inner learning rate of 0.01.
@@
Discretization-invariance is an important property for neural operators. While the output of HyPINO is a continuous PINN that can be evaluated at arbitrary spatial coordinates, the input PDE parametrization (source function and boundary masks/values) is discretized on a fixed-size grid ($224\times 224$) to match the Swin Transformer’s input resolution. Following prior work Herde et al. [2024], this limitation can be mitigated by demonstrating test-time resolution invariance when varying the input grid resolution and resizing it to ($224\times 224$).
@@
Residual-Based Refinement for PINO
To examine whether the residual-based refinement strategy is applicable beyond HyPINO, we applied the same refinement procedure to PINO. Table 4 reports the performance across benchmarks, where superscripts indicate the number of refinement steps.
Table 4: Residual-based refinement applied to PINO across seven PDE benchmarks. Each cell reports mean squared error (MSE). Superscripts denote the number of refinement steps.
HT HZ HZ-G PS-C PS-L PS-G WV
PINO 1.38e-2 2.02e-2 6.09e-2 1.70e-1 3.31e-3 3.05e-1 2.95e-1
PINO3 1.31e-2 7.22e-3 4.61e-2 2.83e-2 4.64e-3 2.33e-2 3.07e-1
PINO10 3.90e-2 5.09e-3 1.44e-1 1.05e-2 1.01e-3 1.76e-2 8.53e-1
We observe consistent improvements with refinement, particularly on HZ, PS-C, and PS-G. The drop in performance on PS-L mirrors the behavior we observed with HyPINO and is likely related to the small target values. These findings suggest that residual-based refinement is applicable to physics-informed neural operators in general, and not limited to our specific architecture.

### v3

@@ 4.3 Evaluation
We evaluate HyPINO and baseline models on seven standard PDE benchmarks from the PINN literature: (i) HT - 1D heat equation [32], (ii) HZ - 2D Helmholtz equation [2], (iii) HZ-G - Helmholtz on an irregular geometry [16], (iv) PS-C - Poisson with four circular interior boundaries [16], (v) PS-L - Poisson on an L-shaped domain [32], (vi) PS-G - Poisson with a Gaussian vorticity field [19], and (vii) WV - 1D wave equation [16]. The exact problem statements and visualizations of the corresponding parameterizations are provided in Appendix B.
Table 1: Model performance across seven PDE benchmarks. Each cell shows mean squared error (MSE) / symmetric mean absolute percentage error (SMAPE) [34]. Lower is better.
@@
Table 1 further highlights the advantages of our proposed iterative refinement approach. After three refinement iterations (HyPINO3), we observe substantial reductions in prediction error across all but one benchmark. Notably, the MSE for PS-C and PS-G decreases by more than one order of magnitude, and an even larger improvement is observed for HT (almost two orders of magnitude). With ten refinement iterations (HyPINO10), our model achieves state-of-the-art performance on all but two evaluated benchmarks, outperforming the best baseline models by factors ranging from 2.1 (on HZ against Poseidon) to 173 (HT against PINO). Table 2 shows that iterative refinement leads to a continuously more accurate prediction on the challenging WV benchmark, with the model being able to extend the undulating shape continuously further from the initial condition across the time dimension. Importantly, our results indicate that iterative refinement is not specific to HyPINO but serves as a generally effective test-time enhancement for other physics-informed neural operators, as demonstrated by the performance of PINO3 and PINO10.
@@ 4.4 Finetuning
The parameters $\theta^{\star}$ produced by HyPINO can be used to initialize PINNs for subsequent fine-tuning on specific PDE instances. We compare the convergence behavior of HyPINO-initialized PINNs with those initialized randomly and with Reptile meta-learning [38], where Reptile was trained on our synthetic dataset using 10,000 outer- and 1,000 inner-loop cycles. We also evaluate ensembles generated with HyPINO3 and HyPINO10 against ensembles of equal size and architecture initialized with random weights or Reptile.
@@ A.1 Neural Operator Architecture
We build on the HyperPINN framework [8] and design a neural operator based on a hypernetwork that generates the weights $\theta^{\star}$ of a Physics-Informed Neural Network (PINN) $u_{\theta}$, conditioned on a given PDE instance. Specifically, the hypernetwork learns a mapping
@@
Each grid-valued input is first passed through a Fourier feature mapping [45], which augments the input with sinusoidal encodings using five exponentially increasing bands $\text{frequencies}=0.1\cdot 2^{i},\quad i\in\{0,1,2,3,4\}$. This enhances the network’s ability to represent high-frequency content and reduces spectral bias. The Fourier mapping layer is followed by two convolutional layers with kernel size three and strides of two. For the boundary location grids $M$ (c.f. Section 3.1, we compute four embeddings: $z_{D}^{1},\;z_{D}^{2},\;z_{N}^{1},\;z_{N}^{2}$. For the boundary value grids $V$, we compute $z_{g}$ (Dirichlet values) and $z_{h}$ (Neumann values). The source term yields the embedding $z_{f}$.
@@
The grid embedding $z_{G}$ is processed by a sequence of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$. Denoting by $z^{(i)}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ the output of block $\mathcal{SW}_{i}$ and $z^{(0)}=z_{G}$, we interleave each block with a FiLM modulation [40] conditioned on the coefficient embeddings $z_{C}$. Concretely, we define
@@
Inspired by Swin Transformer UNet architectures [4, 11], we retain all intermediate latent representations from the Swin blocks $\{z^{(i)}\}_{i=1}^{K}$ to keep information at various semantic levels.
@@
To aggregate spatial information into a compact latent representation suitable for parameterizing the target PINN, we perform Multi-Head Attention Pooling [25, 55] across the flattened outputs of each Swin Transformer block. Specifically, let $z_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ denote the output of the $i$-th FiLM-modulated Swin block. We reshape it into a sequence of tokens $kv_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$, which serve as both keys and values in the attention mechanism.
@@
We define the architecture of the target PINN as MLP with Fourier feature mapping [45] and multiplicative skip connections [46]. Fourier encodings provide spectral expressivity for modeling high-frequency components [47], while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by masking some layers.
@@
Following Wang et al. [46], the encoded input is projected through three parallel transformations:
@@ A.2.5 Sampling Physical Domains
We employ a randomized sampling procedure based on Constructive Solid Geometry (CSG) [33] to generate complex and diverse domains.
@@ B.1 Heat 1D (HT)
This benchmark problem is adapted from DeepXDE [32]. Figure 5 shows the parametrization of the different PDE components.
@@ B.2 Helmholtz 2D (HZ)
This benchmark problem is adapted from DeepXDE [32]. Figure 6 shows the parametrization of the different PDE components.
@@ B.3 Helmholtz 2D - Complex Geometry (HZ-G)
This benchmark problem is adapted from PINNacle [16]. Figure 7 shows the parametrization of the different PDE components.
@@ B.4 Poisson 2D - Circles (PS-C)
This benchmark problem is adapted from PINNacle [16]. Figure 8 shows the parametrization of the different PDE components.
@@ B.5 Poisson 2D - L-Domain (PS-L)
This benchmark problem is adapted from DeepXDE [32]. Figure 9 shows the parametrization of the different PDE components.
@@ B.6 Poisson 2D - Gauss (PS-G)
We select a sample from the dataset introduced in [19]. Figure 10 shows the parametrization of the different PDE components.
@@ B.7 Wave 1D (WV)
This benchmark problem is adapted from PINNacle [16]. Figure 11 shows the parametrization of the different PDE components.
@@ B.8 Baseline Models
UNet [42]. A convolutional encoder–decoder network that shares the same input encoding architecture as HyPINO, but replaces the transformer-based decoder with a purely convolutional upsampling stack that directly outputs a solution grid of shape $(224\times 224)$, matching the resolution of the input tensors. It is trained exclusively on supervised PDEs with analytical solutions, using a batch size of 128, an initial learning rate of $10^{-4}$, and for 30,000 training batches. The UNet has a total parameter count of 62M.
Poseidon [19]. A large pretrained operator network with approximately 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the input dimensionality to 5 to accept all grid-based inputs. Additionally, the lead-time-conditioned layer normalization layers—originally designed to condition on a 1D time input—are modified to condition on the 5D vector of differential operator coefficients. Poseidon is fine-tuned exclusively on supervised samples, using the same training setup as the UNet (30,000 batches, batch size 128, initial learning rate $10^{-4}$).
PINO[29]. A Fourier neural operator[28] architecture with 33M parameters, trained with joint physics-informed and supervised losses computed in Fourier space. We adapt the model to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It follows the same hybrid supervision and training curriculum as HyPINO, including physics-informed losses, and is trained for 30,000 batches with a batch size of 128 and an initial learning rate of $10^{-4}$.
@@ B.10 Finetuning
We investigate the utility of HyPINO-generated PINN parameters $\theta^{\star}$ as a prior for rapid adaptation to specific PDE instances. We compare three initialization strategies: (i) HyPINO-initialized PINNs, (ii) randomly initialized PINNs, and (iii) PINNs initialized via Reptile meta-learning [38]. For Reptile, we use 10,000 outer-loop steps on our synthetic dataset and 1,000 inner-loop updates per sample with an inner learning rate of 0.01.
@@
Discretization-invariance is an important property for neural operators. While the output of HyPINO is a continuous PINN that can be evaluated at arbitrary spatial coordinates, the input PDE parametrization (source function and boundary masks/values) is discretized on a fixed-size grid ($224\times 224$) to match the Swin Transformer’s input resolution. Following prior work [19], this limitation can be mitigated by demonstrating test-time resolution invariance when varying the input grid resolution and resizing it to ($224\times 224$).